Circular No. 6655

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 T. J. Galama and P. J. Groot, University of Amsterdam (UoA); J.

van Paradijs, UoA and University of Alabama in Huntsville; C. Kouveliotou, Universities Space Research Association; M. Centurion, Isaac Newton Group Telescopes (ING) and Instituto Astrofisico de Canarias, J. H. Telting, ING; P. Smith and C. Mackey, Kitt Peak National Observatory (KPNO); and J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, report photometry of the variable star reported by Bond (IAUC 6654), obtained with the 3.5-m WIYN telescope at KPNO on May 9.202 UT and with the 4.2-m William Herschel Telescope on La Palma during May 10.03–10.07 and 11.01–11.05, confirming that this object is brightening: May 9, R = 21.2 ± 0.1; 10, U = 21.5 ± 0.2, V = 21.0 ± 0.1, R = 20.35 ± 0.1, I = 20.9 ± 0.2; 11, U = 20.5 ± 0.2, B = 21.3 ± 0.2, V = 20.2 ± 0.1, R = 20.0 ± 0.1, I = 18.5 (highly uncertain).

S. G. Djorgovski, M. R. Metzger, S. C. Odewahn, R. R. Gal, S. R. Kulkarni, and M. A. Pahre, California Institute of Technology (CIT); D. A. Frail, National Radio Astronomy Observatory; and E. Costa, M. Feroci, and the rest of the BeppoSAX team, write: “We have obtained deep optical images of the field of GRB 970508 at the Palomar 5-m and 1.5-m telescopes on May 9, 10, and 11 UT. Bond’s optical variable is measured at α = 6^{h}53^{m}49^{s}.43, δ = +79^{o}16′19′′.6 (equinox 2000.0), with uncertainties of 1′′. Our preliminary reductions indicate that the source had Gunn r magnitudes of 21.33, 20.17, and 20.15 on May 9.19, 10.23, and 11.21, respectively, corresponding to fluxes f(ν) of 11.1, 32.2, and 32.8 μJy at the effective wavelength of 665 nm (estimated total errors of 10-20 percent). The Gunn colors measured on May 10.23 were g −r = −0.01 and r −i = +0.02, which can be fitted with a power-law slope of ∼ −0.5 in f(ν) vs. ν, with large errors. No other significantly variable sources brighter than r ∼ 23 have been detected in the revised GRB error circle.”

Metzger, Djorgovski, C. C. Steidel (CIT), Kulkarni, K. L. Adelberger (CIT), and Frail further write: “Spectra of Bond’s variable were obtained using the Keck II 10-m telescope (+ LRIS) on May 11.25 UT. The object has no obvious emission lines. We have identified a set of aborption features between 430 and 514 nm that we associate with the Fe II triplet near 237.5 nm, Fe II at 258.6 and 260.2 nm, and Mg II at 279.6 and 280.3 nm, which places this system at z = 0.835. Several other absorption features are also seen. If the source is associated with GRB 970508, it is evident that the γ-ray burst lies at z ≥ 0.835. Absence of prominent Lyα forest in the spectrum also suggest that the source is at redshift z < 2.1.”

1997 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green